Beale Manufacturing Company has a beta of 1.1, and Foley Industries has a beta of 0.5. The required return on an index fund that holds the entire stock market is 11%. The risk-free rate of interest is 7%. By how much does Beale's required return exceed Foley's required return? Round your answer to two decimal places.
Answer #1

Required return=risk free rate+beta*(market rate-risk free rate)

Required return for:

Beale Manufacturing Company=7+(11-7)*1.1=11.4%

Foley Industries=7+(11-7)*0.5=9%

Hence difference=11.4-9

=2.4%